A body of mass m is accelerated uniformly from rest to a speed v in a time t.The instantaneous power deliverd to body as a function of time , is given by

Grade:11

1 Answers

Nirmal Singh.
9 years ago
u = 0
Total time is T
so v = u+aT

v = aT

a = v/T

Hence velocity at any time t, v’ = (v / T)*t

Hence instantaneous power P = F*v’

P = ma*v’

P = m (v/T)*(v/T)*t

P = (mv2/ t2)*t

Regards,

Nirmal SIngh
Askiitians Faculty